OpenStudy (anonymous):

y=3x/x^2+9 list the intercepts. let me guess my y intercept is 0?

OpenStudy (anonymous):

X -> (-1/3, 0) Y -> None

OpenStudy (anonymous):

is this 3x Divided y x^2+9?

OpenStudy (dumbcow):

i disagree y-intercept is (0,0) because y = 3*0/0+9 = 0 x intercept is also (0,0) because 0= 3x/x^2+9, 3*0 = 0

OpenStudy (anonymous):

truee dumbcow the intercept will be when 3x =0

OpenStudy (anonymous):

this is y= 3x divided by x^2+9

OpenStudy (anonymous):

when in doubt use calculator haa

OpenStudy (anonymous):

yes because your intercept for x will be when y = 0

OpenStudy (anonymous):

and the polynomial you are evaluating this will be 0 when the numerator = 0 which is 3x=0 now for y intercepts just solve for when x=0 which will also get you 0

OpenStudy (anonymous):

i agree with pk I don't think the function given is a polynomial it seems more like a rational function

OpenStudy (dumbcow):

it is a rational function

OpenStudy (anonymous):

Exactly!!! \[y = (3x/x ^{2}) + 9\]

OpenStudy (anonymous):

yeah so it doesn't cross the y - axis but intercepts the x - axis at (-1/3,0)

OpenStudy (anonymous):

y-axis is an asymptope

OpenStudy (dumbcow):

\[y = \frac{3x}{x^{2}+9}\]

OpenStudy (dumbcow):

you misread the question ..use parenthesis people

OpenStudy (anonymous):

oh wow okay that's how you see the equation: I solved it as y = 3/x + 9